# Kernel-U-Net: Symmetric and Hierarchical Architecture for Multivariate Time Series Forecasting

### v2

@@ 3.2 Kernel U-Net
Creation of Kernel U-Net
Let us note input length $L$, input dimension $M$, Lists of multiples of look-back window $,[L_{2},$...,… , $L_{n}]$ and feature $,[L_{n+2},$...,… , $L_{n+m}]$ , size of patch $L_{1}$ and feature-unit $M_{1}$ , a list of hidden dimension $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, a list of kernels $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, latent vector length $L_{h}$ and dimension $M_{h}$ Remark that the hidden dimension of the model is the size of intermediate output vectors within layers. It corresponds to channel size in a convolutional network. This dimension can be augmented for a larger passage of information. We describe the creation of the K-U-Net encoder in Algorithm 2 and leave that of the decoder (Algorithm 3) and K-U-Net (Algorithm 4) in the appendix.
3.3 Custom Kernels
Linear kernel
The linear kernel is a simple matrix multiplication (Figure 3). Given $X$ of shape $(B,L_{i},M_{i})$, linear kernel $f_{i}$ reshape it into $(B,L_{i}*M_{i})$ and process it as follow:
$Z_{i}=f_{i}(X)=X*w+b$
, where $Z_{i}$ is output of kernel, $w\in R^{L_{i}M_{i}\times L_{i+1}M_{i+1}}$ is weight matrix and $b\in R^{L_{i+1}M_{i+1}}$ is bias vector. Remark that the number of parameters of $w$ in a linear kernel is $L_{i}*1*ch_{in}*ch_{out}$ which is equivalent to that of a 1D convolutional layer whose kernel size is $L_{i}*1$.
The multi-layer linear kernel has an additional hidden layer and a non-linear activation function Tanh. The formulation is:
$Z_{i}=f_{i}(X)=Tanh(X*w_{1}+b_{1})*w_{2}+b_{2}$
, where $Z_{i}$ is output of kernel, $w_{1}\in R^{L_{i}M_{i}\times L^{\prime}M^{\prime}}$ and $w_{1}\in R^{L^{\prime}M^{\prime}\times L_{i+1}M_{i+1}}$ are weight matrix , $b_{1}\in R^{L^{\prime}M^{\prime}}$ and $b_{2}\in R^{L_{i+1}M_{i+1}}$ are bias vectors, and $L^{\prime}=\frac{1}{2}(L_{i}+L_{i+1}),M^{\prime}=\frac{1}{2}(M_{i}+M_{i+1})$.
Transformer kernel
The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with Relu activation. The transformer kernel in this work follows the classical structure.
LSTM kernel
The LSTM kernel contains a classic LSTM cell and a linear layer for skip connection. The hidden states of all time steps are flattened and used for a linear combination in the next. The detailed structure of the Transformer kernel and the LSTM kernel is discussed in the appendix.
Figure 3: Structure of Linear Kernel. a) Linear kernel, b) Multi-layer Linear Kernel with Tanh activation.
Table 1: Multivariate time series forecasting results with Kernel U-Net. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best results in underlined.
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer Yformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.370 0.393 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713 0.985 0.74
192 0.404 0.414 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792 1.17 0.855
336 0.420 0.430 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809 1.208 0.886
720 0.438 0.454 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865 1.34 0.899
ETTh2
96 0.271 0.335 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525 1.335 0.936
336 0.357 0.4 0.329 0.384 0.357 0.4 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835 1.444 0.96
720 0.39 0.438 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625 3.498 1.631
ETTm1
96 0.286 0.342 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571 0.849 0.669
192 0.330 0.363 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669 0.928 0.724
336 0.360 0.384 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871 1.058 0.786
720 0.405 0.412 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823 0.955 0.703
ETTm2
96 0.16 0.245 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453 0.487 0.529
192 0.215 0.215 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563 0.789 0.705
336 0.268 0.326 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887 1.256 0.904
720 0.343 0.379 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338 2.698 1.297
Electricity
96 0.129 0.226 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368 - -
192 0.147 0.244 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386 - -
336 0.163 0.261 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394 - -
720 0.197 0.292 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439 - -
Traffic
96 0.358 0.253 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391 - -
192 0.373 0.262 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379 - -
336 0.390 0.271 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42 - -
720 0.430 0.292 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472 - -
Weather
96 0.142 0.195 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384 - -
192 0.188 0.244 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544 - -
336 0.241 0.284 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523 - -
720 0.310 0.333 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741 - -
3.4 Complexity Analysis
As the K-U-Net is symmetric, we study the complexity of the encoder. Let us suppose that a kernel is receiving a sequence of length $L$ where $L=\prod_{1}^{n}L_{i}$ and all $L_{i}$ are equal. As the patch size in the first layer of Kernel U-Net encoder is $L_{1}$, a kernel module will process $\frac{L}{L_{1}}$ patches of size $L_{1}$. Therefore, the complexity is $O(\frac{L}{L_{1}^{n}}*f(L_{1}))$ where $f(L_{1})$ is the complexity inside the kernel in the function of the patch size $L_{1}$ and $n$ is the index of the layer. In the case of using the linear kernel at the first layer, the complexity is $O(\frac{L}{L_{1}}*L_{1})$= $O(L)$. In the case of using a classic transformer kernel, the complexity is $O(\frac{L}{L_{1}}*L_{1}^{2})$= $O(L*L_{1})$. Let us set $L_{1}=log(L)$, the complexity of the application of such a quadratic calculation kernel is $O(L*log(L))$. Moreover, if we apply the transformer kernel starting from the second layer, the complexity is dramatically reduced to $O(\frac{L}{L_{1}^{2}}*L_{1}^{2})$= $O(L)$. Following the same demonstration, the complexity of using multi-layer linear and LSTM kernels starting from the second layer is also bounded by $O(L)$.
Figure 4: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on ETTh1 dataset.
4 Experiments and Results
Datasets.
We conducted experiments on our proposed Kernel U-Net on 7 public datasets, including 4 ETT (Electricity Transformer Temperature) datasets [^1] https://github.com/zhouhaoyi/ETDataset (ETTh1, ETTh2, ETTm1, ETTm2), Weather [^2] https://www.bgc-jena.mpg.de/wetter/, Traffic [^3] http://pems.dot.ca.gov and Electricity [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams
20112014
. These datasets have been benchmarked and publicly available onWu et al. (2021) and the description of the dataset is available in Zeng et al. (2023)).
Figure 5: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on Weather dataset.
Here, we followed the experiment setting in Zeng et al. (2023) and partitioned the data into $[12,4,4]$ months for training, validation, and testing respectively for the ETT dataset. The data is split into $[0.7,0.1,0.2]$ for training, validation, and testing for the Weather, Traffic, and Electricity datasets.
Baselines and Experimental Settings.
We follow the experiment setting in NLinear Zeng et al. (2023) and take $L=\{336,720\}$ step historical data as input then forecast $T\in\{96,192,336,720\}$ step value in the future. We replace the last value normalization with mean value normalization for ETT, Electricity, and Weather datasets, and apply instance normalizationUlyanov et al. (2017) for the Traffic dataset. We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as mentioned in Wu et al. (2021).
We include recent methods: PatchTSTNie et al. (2023), NLinear, DLinearZeng et al. (2023), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021), LogTrans Li et al. (2019) YformerMadhusudhanan et al. (2023). We merge the result reported in Nie et al. (2023) and Zeng et al. (2023) for taking their best in a supervised setting and execute Yformer in our environment with default parameters in its Github [^5] https://github.com/18kiran12/Yformer-Time-Series-Forecasting.
Figure 6: Computation time of K-U-Net, PatchTST and Yformer.
Model Variants Search.
We propose 4 kernels for experiments with K-U-Net on 7 datasets. By replacing a linear kernel at different layers with other types of kernels for searching for variants that adapt the dataset. The variants are noted as ”K-U-Net $<$kernel$>$_$<$replacing_layer$>$ ($<$lookback_window$>$)”. For example, K-U-Net Linear_$\_0000(720)$ means that the model is made of linear kernels at all layers and the lookback window size is 720, K-U-Net Transf_$\_0100(720)$ means that a transformer kernel replaces the linear kernel at the second layer, K-U-Net Linear_$\_0110(720)$ means that multilayer linear kernels replace the kernels at the second and third layers. We compose 16 variants with Multi-layer linear kernels and 7 variants with Transformer and LSTM kernels.
$RS(m)=Min_{t\in T}\frac{Top5MMSE(m,t)}{Min_{\{m\in M,l\in L\}}Top5MMSE(m,t,l)}$
, where $m$ is a model in the models set $M$, $t$ is a forecasting horizon in $T$, $l$ is a look back window in $L$. Relative score notes the best-performed model with 1 and thus helps to identify the high-potential candidates for further fine-tuning examinations.
We observe in Figure 4 that the best variant for ETTh1 dataset is K-U-Net Linear_$\_0010$. Comparing the relative score of the K-U-Net of replacing__\__layer code ($0010$, $0011$), ($0110$, $0111$) and ($0100$, $0101$), we remark that replacing the highest layer with Transformer and LSTM kernel degrades the performance because of overfitting. Furthermore, We observe in Figure 5 that the best variants for the Weather dataset are K-U-Net Linear_$\_0110$, K-U-Net LSTM_$\_0100$ and K-U-Net Transf_$\_0100$. Comparing the relative score of the K-U-Net of replacing__\__layer code ($0010$, $0110$), ($0011$, $0111$) and ($0001$, $0101$) we remark that replacing the second layer with Transformer and LSTM kernel gains the performance for their expressiveness.
Among candidates in the search phase, we empirically choose 3 variants for fine-tuning experiments with 5 runs. The final result shows that the performance of Kernel U-Net exceeds or meets the state-of-the-art methods in univariate settings and multivariate settings in most cases.
Multivariate time series forecasting result.
We remark that our model improves the MSE performance around $72.92\%$ compared with Yformer and $0.99\%$ compared to PatchTST and NLinear in the multivariate setting (Table 1). It is worth noting that K-U-Net achieves similar results on the Electricity dataset with a variant based on Multilayer linear kernels.
Univariate time series forecasting result.
Additionally, we remark that our model improves the MSE performance around $60.47\%$ compared with Yformer and $6.06\%$ compared to PatchTST and NLinear in the Univariate setting(Table 3). It is worth noting that K-U-Net improves the MSE performance on ETTh2 for $15.67\%$ compared to PatchTST.
Figure 7: GPU consumation of K-U-Net, PatchTST and Yformer.
Computation Efficiency
We examined the computation efficiency of 4 variants of K-U-Net, PatchTST, and Yformer. We execute the models on the ETTh1 and ETTm1 datasets for 10 epochs and measure the average execution time per epoch and the GPU consumption during the training. The hidden dimension is set to 128 for all models. For fair comparison, PatchTST and Yformer contain 2 layers of Transformer block which equals to K-U-Net_Transf_$\_0100$. All experiments are executed on a Tesla V100 GPU in a Google Colab environment. Comparing to the PatchTST, K-U-Net Linear_$\_0000$ saves $85.20\%$ and $83.07\%$ memory (Figure 7) and saves $81.82\%$ and $80.55\%$ computation time (Figure 6) on ETTh1 and ETTm1 datasets, K-U-Net_Transf_$\_0100$ saves $55.15\%$ and $50.00\%$ memory and saves $28.47\%$ and $26.88\%$ computation time respectively.
5 Conclusion
Appendix A Appendix
A.1 Complement Algorithms
The complete algorithm of Kernel U-Net contains creation of Encoder, Decoder (Algorithm 3), Wrapper, and Kernel U-Net(Algorithm 4).
Algorithm 3 Kernel U-Net Decoder
Input: $L_{h}$, $M_{h}$, $[L_{n},...,L_{2}]$, $[L_{n+m},...,L_{n+2}]$, $L_{1}$, $M_{1}$, $[f_{n+m},...,f_{n+2},$ $f_{n},...,f_{2}]$, $[M_{n+m},...,M_{n+2},$ $M_{n},...,M_{2}]$, $L_{1}$, $M_{1}$.
Output: Instance of Kernel U-Net Decoder
@@ A.1 Complement Algorithms
  def __init(inputs)__:
  # Create a list of multiples
[…]
  # Create first layer
  layers = [KernelWrapper($f_{n+m},$ $1,$ $M_{h},$
  $L_{n+m},$ $M_{n+m}$)]
  # Create intermediate layers
  for  $i$ in $[n+m-1,...,n+2,n,...,2]$ do
           layers.append(KernelWrapper($f_{i},$ $1,M_{i+1}$
           $,L_{i},M_{i})$)
  end for
  # Create last layer
  layers.append(KernelWrapper($f_{2},1,$ $M_{2},$ $L_{1},$ $M_{1})$)
  decoder = nn.Sequential (*layers)
Algorithm 4 Kernel U-Net
Input: $L_{1},M_{1}$, $[,L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$, $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
Output: Instance of Kernel U-Net
  # Define the init function :
  def __init(inputs)__:
  # Initiation of encoder and decoder in K-U-N.
  encoder = KernelUNetEncoder(inputs)
  decoder = KernelUNetDecoder(inputs)
@@
  reshape x to $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$
  transpose x to $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$
  reshape x to $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$
  # Apply encoder on x
  z = encoder(x)
  for  $i$ in $[1,...,n]$ do
           output = encoder.layers[i]._unet_skip_output
           decoder.layers[$n-i+1$]._unet_skip_input
           = output
  end for
  # Apply decoder on z
  y= decoder(z)
  reshape y to $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$
  transpose y to $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$
  reshape y to $(B,L,M)$
  return y
A.2 LSTM and Transformer kernels
Transformer kernel
The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation. We follow the description in Wen et al. [2023].
Positional encoding (PE).
In a vanilla Transformer, the input sequence is added by a positional encoding vector given by:
$PE(t)=\left\{\begin{array}[]{ll}sin(\omega it)&ifi\%2=0\\ cos(\omega it)&ifi\%2=1\\ \end{array}\right.$
, where $\omega i$ is the frequency of waves generated at each dimension. By adding positional information to the input, the transformer can consider the same patterns that occur in different time steps as different patterns.
Figure 8: Structure of Kernels a) Linear Kernel, b) Multi layer linear Kernel with Tanh activation, c) LSTM Kernel, d) Transformer Kernel
Multiple head attentions.
With the Query-Key-Value $(QKV)$ model, the scaled dot-product attention used by Transformer is given by
$Attention(Q,K,V)=Softmax(\frac{QK^{T}}{\sqrt{D_{k}}}))V$
, where queries $Q\in R^{N\times D_{k}}$ , keys $K\in R^{M\times D_{k}}$ , values $V\in R^{M\times D_{v}}$ , $N,M$ denote the lengths of queries and keys (or values), and $D_{k},D_{v}$ denote the dimensions of keys (or queries) and values.
The multiple attention block contains repeated heads and can be given by
$MultiHeadAttn(Q,K,V)=Concat(h_{1},...,h_{H})W^{O}$
, where $h_{i}=Attention(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}).$
The transformer kernel in this work (Figure 8) follows the classical structure with 2 additional linear layers. The first linear layer before the Transformer block of size $M_{in}\times M_{out}$ reduces the dimension. The last linear layer is a matrix of size $L_{in}*M_{out}\times L_{out}*M_{out}$ that reduces the input length.
LSTM kernel
The Long Short-term Memory contains 3 gates: Forget Gate $f_{t}$, Input Gate $i_{t}$, Output Gate $o_{t}$. Each gate has a weight matrix $W$ and a bias vector $b$ for the current input $x_{t}$ and the previous hidden state $h_{t-1}$. The next cell state $C_{t}$ is updated by the current activated cell state $C^{\prime}_{t}$. The next hidden state $h_{t}$ is updated by the output gate and the next activated cell state. The formulation of the LSTM cell is as follows:
$f_{t}=\sigma(W_{f}[h_{t-1},x_{t}]+b_{f})$
$i_{t}=\sigma(W_{i}[h_{t-1},x_{t}]+b_{i})$
$C^{\prime}_{t}=tanh(W_{C}[h_{t-1},x_{t}]+b_{C})$
$C_{t}=f_{t}C_{t-1}+i_{t}C^{\prime}_{t}$
$o_{t}=\sigma(W_{o}[h_{t-1},x_{t}]+b_{o})$
$h_{t}=o_{t}tanh(C_{t})$
In our experiments, we add linear layers (Figure 8) that take the flattened output of all hidden states of the LSTM cell for linear combination.
$Z=W*Flatten([h_{1},...,h_{t}])+b$
, where $W\in R^{L_{i}*d_{h}}$, $b\in R^{L_{i+1}*M_{i+1}}$ and $d_{h}$ is the dimension of hidden state and $d_{h}=M_{i}$ in general. The residual skip operation aside from the LSTM cell contains a linear layer equivalent to a simple linear kernel discussed in the previous sections.
A.3 Dataset and Experiment
Dataset Statistics
We use 7 popular multivariate datasets for experiments. We create subset Weather(S), Electricity(S), and Traffic(S) from top-left to bottom-right for quick search of K-U-Net variants. The statistic is in noted in Table 2. We refer the description of datasets in Nie et al. [2023]:
  • Weather dataset collects 21 meteorological indicators in Germany, such as humidity and air temperature.
  • Traffic dataset records the road occupancy rates from different sensors on San Francisco freeways.
  • Electricity is a dataset that describes 321 customers’ hourly electricity consumption.
  • ETT (Electricity Transformer Temperature) datasets are collected from two different electric transformers labeled with 1 and 2, and each of them contains 2 different resolutions (15 minutes and 1 hour) denoted with m and h.
Table 2: Statistics of popular datasets for experiments.
Datasets ETTh1 ETTh2
ETTm1
ETTm2 Weather
Traffic
Electricity Weather(S) Traffic(S) Electricity(S)
Features 7 7
7
7 21
862
321 21 (100%) 21 (2.5%) 22 (7%)
Timesteps 17420 17420
69680
69680 52696
17544
26304 10539 (20%) 12280 (70%) 10521 (40%)
Table 3: Univariate time series forecasting results with Kernel U-Net. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best in underlined.
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer LogTrans Yformer

### v3

@@ III-A Problem Formulation
III-B Kernel U-Net
Kernel U-Net (K-U-Net) is a neural network featuring hierarchical and symmetric U-shape architecture. It separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels (Figure 1). More precisely, the K-U-Net encoder reshapes the input sequence into a large batch of small patches and repeatedly applies custom kernels on them until the latent vector is obtained. Later, the K-U-Net decoder expands the latent vector into patches gradually at each layer and obtains a large batch of small patches (Figure 1). At last, K-U-Net reshapes the patches to get the final output. In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence and the creation of Kernel-U-Net.
III-B1 Hierarchical partition of the input and output
In the first place, We split the input trajectory matrix into patches. Let us consider a trajectory matrix $X_{t}\in\mathbb{R}^{L\times M}$ at time step $t$. Given a list of multiples $\{L_{2},...,L_{n}\}$ for look-back window and $\{M_{2},...,M_{m}\}$ for feature, the patch length $L_{1}$ and feature unit $M_{1}$ such that $L=\prod_{1}^{n}L_{k}$ and $M=\prod_{1}^{m}M_{k}$. We reshape $X_{t}$ as a set of small patches $P_{t}=\{X_{t,i,j}|X_{t,i,j}\in\mathbb{R}^{L_{1}\times M_{1}}\}$, where $i\in\{1,...,\prod_{2}^{n}L_{k}\}$ and $j\in\{1,...,\prod_{2}^{m}M_{k}\}$. The total number of patches is the product of the multiples of length and feature size #$\#P_{t}=\prod_{2}^{m}M_{k}\cdot\prod_{2}^{n}L_{k}$.
The partitioned patches will be processed by kernels in the encoder gradually and their size will be reduced after the kernel operation at each layer. In the decoder, the patches are generated from vectors of length $1$ at each layer. Since the decoder is symmetrical to the encoder, there will also be a trajectory matrix $\hat{X}_{t+L}\in\mathbb{R}^{T\times M}$ composed of a set of generated patches $\hat{P}_{t}=\{\hat{X}_{t+L,i,j}\}$ as output. For a simpler description, we let the look-back window $L$ and the forecasting horizon $T$ be equal.
Algorithm B Kernel Wrapper (KW)
Input: $\phi$, $X$, $J_{in}$, $D_{in}$, $J_{out}$, $D_{out}$
Output: Instance of Kernel Wrapper
@@ III-B1 Hierarchical partition of the input and output
  def __init__(inputs):
  # Initiation of main kernel function $\phi$
  ϕ=italic-$\phi=$ $\phi(J_{in},D_{in},J_{out},D_{out})$
  # Holds variables for global operation
  skip_out = None
  skip_in = None
@@
  reshape $x$ into (-1, $J_{in}$,$D_{in}$)
  $X=X+$ skip_in if skip_in is not None else $X$
  $Z=\phi(X)$ assert $Z$.shape is (-1, $J_{out}$,$D_{out}$)
  skip_out = $Z$ will be assigned to skip_in in decoder
  return $Z$
III-B2 Hierarchical processing with kernels
The hierarchical processing of Kernel U-Net consists of compressing an input sequence at the encoding stage and generating an output sequence at the decoding stage. By default, kernels reduce the dimension of input at each layer in an encoder and increase the dimension in a decoder. Let us consider $\mathcal{X}\in\mathbb{R}^{B\times L\times M}$, a batch of trajectory matrix $X_{t}$, where $B$ is the batch size. We now describe the shape of the intermediate patch before and after the kernel operation.
At the encoder stage, we compress input $\mathcal{X}\in\mathbb{R}^{B\times L\times M}$ into latent vector $\mathcal{Z}$. Let us assume that $D_{h}$ is the unique dimension of the hidden vectors at each intermediate layer and the latent vector to simplify the problem. Firstly, we reshape $\mathcal{X}$ to $(B,1,\prod_{2}^{n}L_{k},L_{1},\prod_{2}^{m}M_{k},M_{1})$ then transpose it to $(B,\prod_{2}^{m}M_{k},\prod_{2}^{n}L_{k},L_{1},1,M_{1})$ and reshape it to $(B\cdot\prod_{2}^{m}M_{k}\cdot\prod_{2}^{n}L_{k},L_{1},M_{1})$ We denote this vector $\mathcal{P}^{(1)}_{in}$ as a large batch of small patches ready for processing with a kernel. Secondly, the kernel at first layer can now process $\mathcal{P}^{(1)}_{in}$ and outputs a hidden vector $\mathcal{P}^{(1)}_{out}$ in shape $(B\cdot\prod_{2}^{m}M_{k}\cdot\prod_{2}^{n}L_{k},1,D_{h})$. After this operation, we reshape the output $\mathcal{P}^{(1)}_{out}$ to $\mathcal{P}^{(2)}_{in}$ of shape $(B\cdot\prod_{2}^{m}M_{k}\cdot\prod_{3}^{n}L_{k},L_{2},D_{h})$ as input for the next layer. Iteratively, the encoder processes all the multiples $\{L_{2},...,L_{n},M_{2},...,M_{m}\}$ in order, and gives finally a batch of latent vector $\mathcal{Z}=\mathcal{P}^{(n+m)}_{out}$ in shape $(B,1,D_{h})$ (Algorithm A).
TABLE I: Multivariate time series forecasting results with Kernel U-Net. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best results in underlined.
{tblr}
cell11 = c=2, cell13 = c=2, cell15 = c=2, cell17 = c=2, cell19 = c=2, cell111 = c=2, cell113 = c=2, cell115 = c=2, cell117 = c=2, cell21 = c=2, cell31 = r=4, cell71 = r=4, cell111 = r=4, cell151 = r=4, cell191 = r=4, cell231 = r=4, cell271 = r=4, cell381 = c=2, cell383 = c=2, cell385 = c=2, cell387 = c=2, cell389 = c=2, cell3811 = c=2, cell3813 = c=2, cell3815 = c=2, cell391 = c=2, cell411 = r=4, cell451 = r=4, cell491 = r=4, cell531 = r=4, vline2-17 = 1, vline2-3,5,7,9,11,13,15,17 = -, hline1-3,7,11,15,19,23,27,31 = -, rowsep=0.12ex , colsep=3.25ex Methods & K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer Yformer